OpenStudy (anonymous):

how do you solve log base4 (x) + log base4 (x-12)=3

OpenStudy (anonymous):

\[\log_{4} (x)+\log_{4} (x-12)=3 \]

OpenStudy (gorv):

log a*b = log a + log b so log a + log b = log ab \[\log_{4} x +\log_{4} (x-12)=\log_{4} (x*(x-12)=3) \]

OpenStudy (gorv):

\[x*(x-12)=4^{3}=64\]

OpenStudy (gorv):

x^2-12x-64=0 (x-16)(x+4)=0 now do little bit by urself @soccer106